# Emojis Decoded: Leveraging ChatGPT for Enhanced Understanding in Social Media Communications

### v2

@@ 9 Implications and Limitations
Our paper reveals ChatGPT’s precise understanding of emoji functionality and intention. The capability of ChatGPT on emojis could benefit further researchers and social media users. For researchers engaged in future emoji studies, leveraging human-aligned explanations of emoji semantics and intentions could significantly reduce the need for extensive crowd-sourcing efforts, particularly in multilingual contexts. This approach would enable more effective and efficient research into emoji usage and interpretation. As for social media users, they could utilize ChatGPT to decipher the semantics of unfamiliar emojis, thereby enhancing the clarity and transparency of online communication.
@@
There are several limitations in our exploration work. First, previous work has shown that the responses of LLMs are sensitive to temperature and the prompt (Yoo et al. 2021; Schick and Schütze 2020; Gilardi, Alizadeh, and Kubli 2023; Zhou, Maharjan, and Liu 2023), but in our work, for each task, we only explore one type of prompt presented in Appendix. For qualitative and quantitative studies (irony prediction, sentiment classification, and emoji recommendation) in our paper, we set the temperature to 0.7 and 0, respectively.
Second, emoji semantics or sentiment can evolve during the 5-year period and, for example, [Uncaptioned image] (slightly smiling face) has evolved the sense of irony and negative sentiment in Chinese social networks. The human annotations utilized in this paper are almost before 2018, so the 5-year time period may lead to the different annotation results.
@@ 10 Conclusion
.1 Prompt Details
@@ 4. An interesting security vulnerability - albeit not for the everyday car thief
Please assign an intention of emoji usage from the following 8 intentions: sentiment expressed, sentiment strengthened, tone adjusted, content described, content organized, content emphasized, non-communication use, symbol. Please only output the intention answer without justification.
The details of the intentions are suggested as below:

### v3

@@ 9 Implications and Limitations
Our paper reveals ChatGPT’s precise understanding of emoji functionality and intention. ChatGPT’s capabilities with emojis could benefit researchers and social media users. For researchers involved in future emoji studies, leveraging human-aligned explanations of emoji semantics and intentions could significantly reduce the need for extensive crowd-sourcing, particularly in multilingual contexts. This approach would enable more effective and efficient research into emoji usage and interpretation. Social media users could utilize ChatGPT to decipher the semantics of unfamiliar emojis, thereby enhancing the clarity and transparency of online communication.
In addition to ChatGPT’s overall effectiveness in emoji-related tasks, we observed instances of hallucination. For example, when explaining the semantics of the [Uncaptioned image] emoji in Figure 2, emojis with unique meanings on Chinese social media, or when identifying frequently used emojis by French users in Table 4, ChatGPT often relies on spurious correlations or prototypes formed during training or alignment, leading to hallucinations. These findings are valuable for future emoji research, reminding researchers to exercise caution regarding ChatGPT’s hallucinations, especially in cross-cultural contexts. They also highlight the need for model developers to focus more on improving non-verbal token understanding in cross-cultural settings.
@@
With the rapid evolution of current large language models (LLMs), there may be concerns about whether the observations and conclusions presented in this study will still hold for future versions of ChatGPT. On the one hand, in most scenarios where ChatGPT produces human-aligned responses, we anticipate that its accuracy will improve as the model continues to evolve. On the other hand, we also identified cases where ChatGPT produces hallucinations, particularly in the cross-cultural contexts discussed in Sections 6 and 3.3. Since knowledge bases related to platforms like WeChat and Twitter are either not publicly available or restricted for commercial use, we infer that these hallucinations may arise from a lack of relevant knowledge in the training data. We argue that these issues may not be easily resolved by simply scaling up data or model size. Consequently, we expect that the current limitations we have observed are likely to persist and may not be quickly addressed. Therefore, our findings, which delineate the boundaries of ChatGPT’s capabilities, may help researchers better utilize the tool for emoji annotation tasks over time.
There are several limitations in our exploration work. First, previous work has shown that the responses of LLMs are sensitive to temperature and the prompt (Yoo et al. 2021; Schick and Schütze 2020; Gilardi, Alizadeh, and Kubli 2023; Zhou, Maharjan, and Liu 2023), but in our work, for each task, we only explore one type of prompt presented in Appendix. For the quantitative studies in our paper, such as intention prediction, irony prediction, sentiment classification, and emoji recommendation, we use the ChatGPT API (gpt-4-1106-preview and gpt-3.5-turbo-1106) with a temperature setting of 0 to generate the model outputs. For the qualitative studies (emoji semantics and emoji usage within communities), we initially used the ChatGPT web interface to obtain results. To demonstrate the robustness of our findings, we repeated qualitative studies using the ChatGPT API with a temperature of 0 and presented the results in the Appendix .1. Additionally, to gain further insights into open-source models, we replicated some experiments using Llama3-8B with a temperature setting of 0, and the results are also provided in the Appendix .2. We found that the open-source Llama3-8B model exhibits significant hallucinations when probing emoji intentions or probing the associations between emoji and community information, making it unreliable for replacing human annotators with the base Llama3-8B model.
Second, emoji semantics or sentiment can evolve over time. For example, [Uncaptioned image] (slightly smiling face) has evolved the sense of irony and negative sentiment in Chinese social networks. The human annotations utilized in this paper are almost all prepared before 2018, and the 5-year gap may lead to the different annotation results.
@@ 10 Conclusion
.1 Qualitative Studies with ChatGPT API
Emoji Semantics
We repeated the experiments in Section 3 using the ChatGPT API and found that nearly all responses were consistent with those presented in Figures 1 and 2. For example, in the last example in Figure 2, the ChatGPT API continues to interpret the [Uncaptioned image] emoji as representing untrustworthy people rather than as an animal. Furthermore, for the emojis [Uncaptioned image], [Uncaptioned image], [Uncaptioned image], and [Uncaptioned image] in Section 3.3, [Uncaptioned image] and [Uncaptioned image] received correct interpretations when prompted in Chinese, while [Uncaptioned image] and [Uncaptioned image] returned the expected responses when ChatGPT was asked to consider Chinese pop culture. This aligns with the patterns observed in Section 3.3.
Emoji Usage Associated with Communities
We also repeated the experiments from Section 6 using the ChatGPT API, and the results are presented in Tables 9, 10, and 11. The model’s responses from the ChatGPT API with a temperature of 0 closely align with the results obtained via the web interface in Section 6. Therefore, the conclusions that ChatGPT relies on stereotypical associations to infer emoji usage patterns across different communities, as discussed in Section 6, remain valid.
User Emoji from actual usage Emoji from ChatGPT
French user
(Lu et al. 2016)
Female user
(Chen et al. 2018)
Male user
(Chen et al. 2018)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 9: Outputs from ChatGPT API (gpt4-1106-preview) and emojis from actual usage (Lu et al. 2016; Chen et al. 2018) associated with French users, female users, and male users.
Hashtag Emoji from actual usage Emoji from ChatGPT
Table 10: Outputs from ChatGPT API (gpt4-1106-preview) and emojis from actual usage (Zhou and Ai 2022) associated with different hashtags.
Source Emojis associated with GitHub
Table 11: Outputs from ChatGPT API (gpt4-1106-preview) and emojis from actual usage (Zhou et al. 2023a) associated with the GitHub platform.
.2 Llama3 Experiments
To probe the open-source large language models’ capabilities on emoji-related tasks, we choose the state-of-the-art (SoTA) models Llama3-8B (Touvron et al. 2023) to replicate the experiments in Section 3, 4, 5 and 6.
Emoji Semantics
We first ask Llama3 to explain the emoji semantics without context as Section 3.1 but observe that it usually generates hallucinations for the outputs. For example, for the cases in Figure 1, Llama3 generates the correct explanation for [Uncaptioned image], where the detailed output is: I think it is a shrug emoji. It is used to express that you don’t know or don’t care about something. However, for [Uncaptioned image] and [Uncaptioned image], Llama3 directly outputs: I’m not sure what it means and we do not get the desired answers for emoji semantics.
As for the emoji semantics with context, Llama3 does not generate the correct explanations. For example, for the [Uncaptioned image] and [Uncaptioned image] emojis in Figure 2, Llama3 will directly state that The emoji is used to represent the bee or snake, rather than analyzing the meaning of the emoji according to the context.
Emoji Sentiment
For emoji sentiment, we duplicate the experiment for annotating the sentiment of the same 50 emojis in Section 4 and find that Llama3 labels the same sentiment as human annotators in 41/50 cases, which has a lower accuracy than the ChatGPT annotation (49/50).
Emoji Intention
We also repeated the experiments in Section 5.1 using the same prompt to ask the Llama3 model to annotate emoji intentions. However, we found that Llama3 only produced a single token: “\” and then ended the generation. This indicates that Llama3 struggles to follow long instructions with intention rubrics, resulting in meaningless output.
Emoji Usage Associated with Communities
To probe whether Llama3 encodes the knowledge of associations between emojis and communities, we utilize the same prompt in Section 6 and ask Llama3 to generate 10 emojis associated with different communities. However, we find that, in addition to the hashtag #TheLastDance, where it generates a basketball emoji [Uncaptioned image], it produces meaningless hallucinations, such as the code to calculate associations.
Overall, Llama3-8B’s performance falls short compared to human annotators in nearly all emoji-related tasks. The reasons for its shortcomings can be attributed to two main factors. First, as observed in its explanations of emoji sentiment and semantics, Llama3 lacks the necessary knowledge of emojis. Second, due to its limited instruction-following capabilities, Llama3 often generates hallucinations when faced with long or complex prompts. Consequently, we conclude that Llama3-8B is not a suitable replacement for human annotators nor a reliable tool for answering users’ questions about emojis.
.3 Prompt Details
@@ 4. An interesting security vulnerability - albeit not for the everyday car thief
Please assign an intention of emoji usage from the following 8 intentions: sentiment expressed, sentiment strengthened, tone adjusted, content described, content organized, content emphasized, non-communication use, symbol. Please only output the intention answer without justification. The details of the intentions are suggested as below: